OpenStudy (anonymous):

How do I do this? Volume (L) Pressure (atm) PV Product (P × V) 7.3358957 1.0000000 2.4452984 3.0000000 1.2226492 6.0000000 8.150995e-1 9.0000000

OpenStudy (anonymous):

What is the original question?

OpenStudy (anonymous):

9. Complete the tables from the data saved in your lab book. Use only a sampling of the data for pressures at 1, 3, 6, and 9 atm. Ideal Gas 1 MW = 4 g/mol Volume (L) Pressure (atm) PV Product (P × V) 7.3358957 1 2.4452985 3 1.2226492 6 8.150995e-1 9 Ideal Gas 8 MW = 222 g/mol Volume (L) Pressure (atm) PV Product (P × V) 7.3358957 1.0000000 2.4452984 3.0000000 1.2226492 6.0000000 8.150995e-1 9.0000000 Real Gas N2 Volume (L) Pressure (atm) PV Product (P × V) 813.8 1.0000000 2.4438061 3.0000000 1.2212500 6.0000000 8137625e-1 8.9999999

OpenStudy (aaronq):

you multiply the two values together Pressure X Volume

OpenStudy (anonymous):

so its just 7.3358957 X 1=7.3358957 ? I was putting way to much thought into it

OpenStudy (aaronq):

yep thats right, you gotta do it for every row though. I think they just want you to see how these values would differ between an ideal gas and a real gas.

OpenStudy (anonymous):

I was thinking it was going to be much more complicated. I tend to way over think things. But simply multiplying seemed way to easy

OpenStudy (aaronq):

yeah, you shouldn't overcomplicate things. this is pretty straight forward they want the PV product. "product" means the result of a multiplication

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Can you explain this one to me Data Table mass NaCl 4.0801 mass water 99.8g/mL boiling temp of pure water 100.41 c boiling temp of solution 101.15 c 7. What is The boiling point elevation can be predicted using the equation ∆T = Kb x m x i , where ∆T is the change in boiling point, i is the number of ions in the solution per mole of dissolved NaCl (i = 2), m is the molality of the solution, and Kb is the molal boiling point constant for water which is 0.51°C/m. Calculate the predicted change in boiling point, in EC for your solution.

OpenStudy (anonymous):

I haven't had a science class in 25 years and never had a chemistry and this is all Greek to me

OpenStudy (aaronq):

The BP elevation is a colligative property that arises from the presence of a non-volatile solute in a solution, regardless of what the solute is. This is a result of solute particles at the surface interface impeding solvent molecules from entering the gas phase (basically blocking them). In the equation, \(\Delta T\) is the change in temperature from the BP of the pure solvent to that of the solution, \(i\) is the van't hoff constant, \(m\) is molality, and Kb is a constant for a specific solvent. The first step is to find the molality. Do you know the formula for that?

OpenStudy (anonymous):

No,

OpenStudy (aaronq):

okay, so it's pretty simple, \(molality=\dfrac{moles~of~solute}{kilograms~of~solution}\) ------------------------------------------------------------ I'm assuming you know how to find the moles of solute, \(moles=\dfrac{mass}{Molar~ mass}\) Molar mass is calculated from the formula, for NaCl, we add up the molar mass of sodium and the molar mass of chlorine (these are found on the periodic table). Molar mass of sodium + molar mass of chlorine = 22.98 g/mol +35.5 g/mol =58.44 g/mol so, \(moles=\dfrac{4.0801 ~g}{58.44~g/mol}\) ----------------------------------------

OpenStudy (anonymous):

ok and then you divide that? right?

OpenStudy (aaronq):

yes, after you used those moles with the molality equation. Note that the mass of water you were given was in grams and you need to use the value in kilograms (1 kg=1000g).

OpenStudy (aaronq):

i meant "yes, after you use* those moles with the molality equation.", not in the past tense.

OpenStudy (aaronq):

I have to go, hopefully someone else can help you with the rest

OpenStudy (anonymous):

ok thank you so much I'm not quite as lost as I was